# AniSora: Exploring the Frontiers of Animation Video Generation in the Sora Era

### v3

@@ Abstract
Animation has gained significant interest in the recent film and TV industry. Despite the success of advanced video generation models like Sora, Kling, and CogVideoX in generating natural videos, they lack the same effectiveness in handling animation videos. Evaluating animation video generation is also a great challenge due to its unique artist styles, violating the laws of physics and exaggerated motions. In this paper, we present a comprehensive system, AniSora, designed for animation video generation, which includes a data processing pipeline, a controllable generation model, and an evaluation dataset. Supported by the data processing pipeline with over 10M high-quality data, the generation model incorporates a spatiotemporal mask module to facilitate key animation production functions such as image-to-video generation, frame interpolation, and localized image-guided animation. We also collect an evaluation benchmark of 948 various animation videos, the evaluation on VBench and human double-blind test demonstrates consistency in character and motion, achieving state-of-the-art results in animation video generation. Our evaluation benchmark will be publicly available at https://github.com/bilibili/Index-anisora.
Figure 1: Overview. We propose AniSora, a comprehensive framework for animation video generation that integrates a high-quality animation dataset, a spatiotemporal conditional model, and a specialized animation video benchmark. The Data Processing Pipeline constructs a 10M video clip dataset derived from 1M diverse long animation videos. The Video Generation model employs a spatiotemporal conditional model, supporting various User Control and Interaction modes and enabling tasks such as frame interpolation, localized guidance, and so on. The benchmark set comprises 948 ground-truth videos spanning diverse styles, common motions, and both 2D and 3D animations. The prompt suite provides standardized prompts and guiding conditions, complemented by Human Preference Evaluation and a Quantitative Evaluation with eight objective metrics for assessing visual appearance and consistency. AniSora surpasses SOTA models, establishing a new benchmark for animation video generation.
@@ 1 Introduction
footnotetext: *Equal contributions. Core contributors: {jiangyudong, xubaohan, yangsiqian, yinmingyu, liujing}@bilibili.com
The animation industry has seen significant growth in recent years, expanding its influence across entertainment, education, and even marketing. As demand for animation content rises, the need for efficient production processes is also growing quickly, particularly in animation workflows. Traditionally, creating high-quality animation has required extensive manual effort for tasks like creating storyboards, generating keyframes, and inbetweening, making the process labor-intensive and time-consuming. Previous efforts [21, 28] to incorporate computer vision techniques have assisted animators in generating inbetween frames for animation. However, these methods often show effectiveness only within certain artistic styles, limiting their applicability to the varied demands of modern animations.
With recent advancements in video generation, there has been notable progress in generating high-quality videos across various domains. Inspired by Generative Adversarial Networks [9], Variational Autoencoders [13], and, more recently, transformer-based architectures [22, 17], the field has seen remarkable improvements in both efficiency and output quality. However, most video generation methods are trained and evaluated on general-purpose datasets, typically featuring natural scenes or real-world objects [3, 29]. The domain of animation video generation, which plays an important role ranging from entertainment to education, has received relatively little attention. Animation videos often rely on non-photorealistic elements, exaggerated expressions, and non-realistic motion, presenting unique challenges that current methods do not address.
In addition to the generation challenges, the evaluation of video generation is also inherently complex. Evaluating video generation quality requires assessing not only the visual fidelity of each frame but also temporal consistency, coherence, and smoothness across frames [11]. This challenge intensifies in animation, where unique artistic styles must remain consistent despite exaggerated motions and transformations. Progress in this field demands effective evaluation datasets tailored to animated video generation, enabling comprehensive testing of model adaptability to diverse styles, scene changes, and complex motions, thereby driving model optimization and innovation.
@@
In this paper, as shown in Fig. 1, a full system AniSora is presented for animation video generation. First, our data processing pipeline offers over 10 million high-quality text-video pairs, forming the foundation of our work. Secondly, we develop a unified diffusion framework adapted for animation video generation. Our framework leverages spatiotemporal masking to support a range of tasks, including image-to-video generation, keyframe interpolation, and localized image-guided animation. By integrating these functions, our system bridges the gap between keyframes to create smooth transitions and enables dynamic control over specific regions, such as animating different characters speaking precisely. This enables a more efficient creative process for both professional and amateur animation creators. Fig. 2 demonstrates some examples generated by our model under image-to-video conditions.
Additionally, we propose a benchmark dataset specifically designed for animation video evaluation. Unlike existing evaluation datasets, which primarily focus on natural landscapes or real-world human actions, our dataset addresses the unique requirements of animation video assessment. To achieve this, we collected 948 animation videos across various categories and manually refined the prompts associated with each video.
@@
  • We develop a comprehensive video processing system that significantly enhances preprocessing for video generation.
@@
  • We release a benchmark dataset specifically for evaluating animation video generation.
@@ 2.1 Video generation models
With the development of diffusion models, significant progress has been made in video generation over the past two years. Some research including [3, 29, 32, 14] have demonstrated promising results in general video generation. Due to the limited available animation datasets, these models are not particularly effective for animation video generation.
@@ 2.2 Animation video datasets
Video data is one of the most critical elements for generation models, particularly for domain-specific data. However, obtaining high-quality animation video data is especially difficult compared to natural video datasets. Previous research has released some animation-related datasets, including ATD-12K [21], AVC [27]. While these datasets, collected from various animation movies, are helpful for some video interpolation and super-resolution tasks, they are limited by their small size. More recently, Sakuga-42M [33] has been proposed with 1.2M clips. It has improved compared to previous datasets that only contained a few hundred clips. Nevertheless, this remains insufficient for training video generation models, in contrast to general video datasets like Panda-70M [5] and InternVid-200M [26]. Additionally, 80% of its clips are low-resolution and less than 2 seconds, which hampers the generation of high-quality videos.
@@ 2.3 Evaluation of video generation models
Evaluating video generation models has remained a significant challenge in the past few years. Recently, Liu et al. have made great efforts to generate a diverse and comprehensive list of 700 prompts using LLM [15]. Besides, Huang et al. have proposed vbench for general video generation [11]. The authors have released 16 evaluation dimensions and prompt suites. Moreover, there is a notable absence of dedicated animation evaluation datasets, which limits the ability to benchmark models specifically designed for this genre. In [31], the authors primarily have focused on the performance of recent video generation models across various categories of datasets. Furthermore, they have also investigated some vertical-domain models like pose controllable human generation, and audio-driven animation.
@@ 3 Dataset
We build our animation dataset according to the observation that high quality text-video pairs are the cornerstone of video generation, which is proved by recent researches [18]. In this section, we give a detailed description of the construction of our animation dataset and the evaluation benchmark.
Animation Dataset Construction: We build a pipeline to get high-quality text-video pairs among $1$ million raw animation videos. First of all, we use scene detection [4] to divide raw animation videos into clips. Then, for each video clip, we construct a filter rule from four dimensions: text-cover region, optical flow score, aesthetic score, and number of frames. The filter rule is gradually built up through the observations in model training. In detail, the text-cover region score (obtained by [2]) can drop those clips with text overlay similar to end credits. Optical flow score [19] prevents those clips with still images or quick flashback scenes. Aesthetic score [6] is utilized to preserve clips with high artistic quality. Besides, we retain the video clips whose duration is among $2s$-$20s$ according to the number of the frames. After the four steps mentioned above, about $10\%$ clips (more than 10 million clips) can be retained into training step. In addition, a few higher quality clips will be finally filtered from training set to further improve the model’s performance. Specifically, during the training process, we adjust the proportions of specific training data (e.g., talking and motion amplitude) according to the observed performance.
Benchmark Dataset Construction: Moreover, to compare the generation videos between our model and other recent researches directly, we construct a benchmark dataset manually. $948$ animation video clips are collected and labeled with different actions, e.g., talking, walking & running, eating, kissing, and so on. Among them, there are $857$ 2D animation clips and $91$ 3D clips. These action labels are summarized from more than $100$ common actions with human annotation. Each label contains $10$-$30$ video clips. The corresponding text prompt is generated by Qwen-VL2 [25] at first, then is corrected manually to guarantee the text-video alignment.
@@ 4 Method
In this section, we present an effective approach for animation video generation using a diffusion transformer architecture. Section 4.1 provides an overview of the foundational video diffusion transformer model. In section 4.2, we introduce a spatiotemporal mask module that extends the diffusion transformer model, enabling crucial animation production functions such as image-to-video generation, frame interpolation, and localized image-guided animation within a unified framework. These enhancements are essential for professional animation production. Finally, section 4.3 details the supervised fine-tuning strategy employed on the animation dataset.
@@ 4.1 Dit-based Video Generation Model
We adopt a DiT-based [17] text-to-video diffusion model as the foundation model. As shown in Fig. 3, the model leverages the three components to achieve coherent, high-resolution videos aligned with text prompts.
Figure 3: Method. This figure illustrates the Masked Diffusion Transformer framework for animation video generation, designed to support various spatiotemporal conditioning methods for precise and flexible animation control. A 3D Causal VAE compresses spatial-temporal features into a latent representation, generating the guide feature sequence $G$, while a reprojection network constructs the mask sequence $M$. These components, combined with noise and prompt’s feature, serve as input to the Diffusion Transformer. The transformer employs techniques such as patchify, 3D-RoPE embeddings, and 3D full attention to effectively capture and model complex spatial-temporal dependencies. This framework enables seamless integration of features like keyframe interpolation, motion control, and mid-frame extension, streamlining animation production and enhancing creative possibilities.
3D Casual VAE used in video generation frameworks [10, 30]serves as a specialized encoder-decoder architecture tailored for spatiotemporal data compression. This 3D VAE compresses videos across both spatial and temporal dimensions, significantly reducing the diffusion model computing. We follow the approach of Yang et al. [29] to extract latent features, transforming the original video with dimensions $(W,H,T,3)$ into a latent representation of shape $(W/8,H/8,T/4,16)$.
Patchify is a critical step for adapting vision tasks to transformer-based architectures [1]. Given an input video of size $T\times H\times W\times C$, it is split spatio into patches of size $P\times P$, and temporal into size $\ Q$ resulting in $(T/Q)\times(H/P)\times(W/P)\times C$ patches. This method enables efficient high-dimensional data processing by reducing complexity while retaining local spatial information.
3D Full Attention is a module we propose for spatial and temporal modeling, inspired by the remarkable success of long-context training in large language models (LLMs) [8] and foundation video generation models [29, 18].
@@
To stabilize training, we use the v-prediction loss [20], where $v=\sqrt{1-\alpha_{t}}\,x_{0}-\sqrt{\alpha_{t}}\,\epsilon$ and the loss becomes
@@ 4.2 Spatiotemporal Condition Model
Motion Control, as a technique within our framework, addresses the limitations of text-based control and enables precise control over motion regions. This approach enhances artists’ control over video content, allowing them to express their creativity while significantly reducing their workload.
@@ 4.2.2 Motion Area Condition
This framework can also support spatial motion area conditions inspired by Dai et.al [7]. Given the image condition $F_{p_{i}}$, and motion area condition is represented by mask $M_{F}$, the same shape with $F_{p_{i}}$. Motion area in $M_{F}$ is labeled 1, other place is set to 0. As equation 1 in 4.2.1, for guide frame position $p_{i}$, set $M_{p_{i}}=M_{F}$. The data processing and training pipeline can be summarized as follows: Constructing video-mask pairs, we first construct paired training data consisting of videos and their corresponding masks. Using a foreground detector by Kim et.al [12], we detect the foreground region in the first frame of the video. This region is then tracked across subsequent frames to generate a foreground mask for each frame. Union of foreground masks, the per-frame foreground masks are combined to create a unified mask $M_{F}$, representing the union of all foreground regions across the video. Video latent post-processing, for the video latent representation $z_{0}$, non-moving regions are set to the latent features of the guide image, ensuring static areas adhere to the guide. LoRA-based conditional training, we train the conditional guidance model using Low-Rank Adaptation (LoRA) with a parameter size of 0.27B. This approach significantly reduces computational requirements while enabling efficient model training.
@@ 4.3 Supervised Fine-Tuning
Weak to Strong. Our video generation model adopts a weak-to-strong training strategy to progressively enhance its learning capabilities across varying resolutions and frame rates. Initially, the model is trained on 480P videos at 8fps for 3 epochs, allowing it to capture basic spatiotemporal dynamics at a lower frame rate. Following this, the model undergoes training on 480P videos at 16fps for an additional 1.9 epochs, enabling it to refine its temporal consistency and adapt to higher frame rates. Finally, the model is fine-tuned on 720P videos at 16fps for 2.3 epochs, leveraging the previously learned features to generate high-resolution, temporally coherent video outputs. Additionally, we applied stricter filtering as in section3, producing a 1M ultra high-quality dataset for final-stage fine-tuning, significantly boosting high-resolution video quality.
Removing Generated Subtitles. The presence of a significant number of videos with subtitles and platform watermarks in our training data led to the model occasionally generating such artifacts in its outputs. To mitigate this issue, we performed supervised fine-tuning using a curated dataset of videos entirely free of subtitles and watermarks. This dataset, consisting of 790k video clips, was constructed through proportional cropping of videos containing subtitles and the selection of clean, subtitle-free videos. Full-parameter fine-tuning was then applied to the model, and after 5.5k iterations, we observed that the model effectively eliminated the generation of subtitles and watermarks without compromising its overall performance.
Temporal Multi-Resolution Training. Given the scarcity of high-quality animation data, we employ a mixed training strategy using video clips of varying durations to maximize data utilization. Specifically, a variable-length training approach is adopted, with training video durations ranging from 2 to 8 seconds. This strategy enables our model to generate 720p video clips with flexible lengths between 2 and 8 seconds.
@@
Multi-Task Learning. Compared to the physically consistent motion patterns in the real world, animation styles, and motion dynamics can vary significantly across different works. This domain gap between datasets often leads to substantial quality differences in videos generated from guide frames with different artistic styles. We incorporate image generation into a multi-task training framework to improve the model’s generalization across diverse art styles. Experimental results demonstrate that this approach effectively reduces the quality gap in video generation caused by stylistic differences in guide frames.
@@
Mask Strategy. During training, we unmask the first, last, and other frames obtained through uniform sampling with a 50% probability. This strategy equips the model with the ability to handle arbitrary guidance, enabling it to perform tasks such as in-betweening, first-frame continuation, and arbitrary frame guidance, as discussed in Section 4.2.1.

### v4

@@ Abstract
Animation has gained significant interest in the recent film and TV industry. Despite the success of advanced video generation models like Sora, Kling, and CogVideoX in generating natural videos, they lack the same effectiveness in handling animation videos. Evaluating animation video generation is also a great challenge due to its unique artist styles, violating the laws of physics and exaggerated motions. In this paper, we present a comprehensive system, AniSora, designed for animation video generation, which includes a data processing pipeline, a controllable generation model, and an evaluation benchmark. Supported by the data processing pipeline with over 10M high-quality data, the generation model incorporates a spatiotemporal mask module to facilitate key animation production functions such as image-to-video generation, frame interpolation, and localized image-guided animation. We also collect an evaluation benchmark of 948 various animation videos, with specifically developed metrics for animation video generation. Our entire project is publicly available on https://github.com/bilibili/Index-anisora/tree/main
Figure 1: Overview. We propose AniSora, a comprehensive framework for animation video generation that integrates a high-quality animation dataset, a spatiotemporal conditional model, and a specialized animation video benchmark. The Data Processing Pipeline constructs a 10M video clip dataset derived from 1M diverse long animation videos. The Video Generation model employs a spatiotemporal conditional model, supporting various User Control and Interaction modes and enabling tasks such as frame interpolation, localized guidance, and so on. The benchmark set comprises 948 ground-truth videos spanning diverse styles, common motions, and both 2D and 3D animations. The prompt suite provides standardized prompts and guiding conditions, complemented by a Quantitative Evaluation with six objective metrics for assessing visual appearance and consistency. Additionally, Human Preference Evaluation confirms strong alignment with the proposed metrics. AniSora surpasses SOTA models, establishing a new benchmark for animation video generation.
@@ 1 Introduction
00footnotetext: *Equal contributions. Core contributors: {jiangyudong, xubaohan, yangsiqian, yinmingyu, liujing}@bilibili.com
The animation industry has seen significant growth in recent years, expanding its influence across entertainment, education, and even marketing. As demand for animation content rises, the need for efficient production processes is also growing quickly, particularly in animation workflows. Traditionally, creating high-quality animation has required extensive manual effort for tasks like creating storyboards, generating keyframes, and inbetweening, making the process labor-intensive and time-consuming. Previous effortsSiyao et al. (2021); Xing et al. (2024) to incorporate computer vision techniques have assisted animators in generating inbetween frames for animation. However, these methods often show effectiveness only within certain artistic styles, limiting their applicability to the varied demands of modern animations.
With recent advancements in video generation, there has been notable progress in generating high-quality videos across various domains. Inspired by Generative Adversarial NetworksGoodfellow et al. (2014), Variational AutoencodersKingma (2013), and, more recently, transformer-based architecturesVaswani (2017); Peebles and Xie (2023), the field has seen remarkable improvements in both efficiency and output quality. However, most video generation methods are trained and evaluated on general-purpose datasets, typically featuring natural scenes or real-world objectsBlattmann et al. (2023); Yang et al. (2024). The domain of animation video generation, which plays an important role ranging from entertainment to education, has received relatively little attention. Animation videos often rely on non-photorealistic elements, exaggerated expressions, and non-realistic motion, presenting unique challenges that current methods do not address.
In addition to the generation challenges, the evaluation of video synthesis is also inherently complex. Evaluating video generation quality requires assessing not only the visual fidelity of each frame but also temporal consistency, coherence, and smoothness across framesHuang et al. (2024). For animation video generation, this challenge is amplified. Animation videos feature unique artist styles, including color and style that need to be consistent, even as characters undergo exaggerated motions and transformations. Traditional evaluation metrics are commonly used for real-world videos, which may not fully capture the consistency of the main characters and art style of this kind of video. Therefore, developing effective evaluation datasets and metrics customized for animation video generation is essential in this specialized field.
@@
In this paper, as shown in Fig. 1, a full system AniSora is presented for animation video generation. First, our data processing pipeline offers over 10 million high-quality text-video pairs, forming the foundation of our work. Secondly, we develop a unified diffusion framework adapted for animation video generation. Our framework leverages spatiotemporal masking to support a range of tasks, including image-to-video generation, keyframe interpolation, and localized image-guided animation. By integrating these functions, our system bridges the gap between keyframes to create smooth transitions and enables dynamic control over specific regions, such as animating different characters speaking precisely. This allows a more efficient creative process for both professional and amateur animation creators. Fig. 2 demonstrates some examples generated by our model under image-to-video conditions.
Additionally, we propose a benchmark dataset and evaluation metrics specifically designed for animation video evaluation. Currently, there is no existing evaluation dataset for this purpose, so we collected 948 animation videos across various categories and manually refined the prompts. Besides, existing evaluation standards struggle to effectively assess the quality of animation video generation. To address this gap, we have introduced innovative metrics that specifically assess animation video generation. These include character consistency, animation art style consistency, and distortion detection, which are crucial for maintaining the unique visual identity of animation videos.
@@
  • We develop a comprehensive video processing system that significantly enhances preprocessing for animation video generation.
@@
  • To the best of our knowledge, we for the first time released a benchmark dataset and evaluation metrics specifically for evaluating animation video generation.
@@ 2.1 Video generation models
With the development of diffusion models, significant progress has been made in video generation over the past two years. Some research includingBlattmann et al. (2023); Yang et al. (2024); Zheng et al. (2024); Lab and etc. (2024) have demonstrated promising results in general video generation. Due to the limited available animation datasets, these models are not particularly effective for animation video generation.
@@ 2.2 Animation video datasets
Video data is one of the most critical elements for generation models, particularly for domain-specific data. However, obtaining high-quality animation video data is especially difficult compared to natural video datasets. Previous research has released some animation-related datasets, including ATD-12KSiyao et al. (2021), AVCWu et al. (2022). While these datasets, collected from various animation movies, are helpful for interpolation and super-resolution tasks, they are limited by small size. More recently, Sakuga-42MZhenglin Pan (2024) has been proposed with 1.2M clips. It has improved compared to previous datasets that only contained a few hundred clips. However, this remains insufficient for training video generation models, in contrast to general video data sets such as Panda-70MChen et al. (2024) and InternVid-200MWang et al. (2023). Additionally, 80% of its clips are low-resolution and less than 2 seconds, which hampers the generation of high-quality videos.
@@ 2.3 Evaluation of video generation models
Evaluating video generation models has remained a significant challenge in the past few years. Recently, Liu et al. have made great efforts to generate a diverse and comprehensive list of 700 prompts using LLMLiu et al. (2023). Besides, Huang et al. have proposed vbench for general video generationHuang et al. (2024). The authors have released 16 evaluation dimensions and prompt suites. While these dimensions are still insufficient to comprehensively evaluate all aspects of animation video generation. Moreover, there is a notable absence of dedicated animation evaluation datasets, which limits the ability to benchmark models specifically designed for this genre. In Zeng et al. (2024), the authors have focused mainly on the performance of recent video generation models in various categories of datasets. Furthermore, they have also investigated some vertical-domain models like pose-controllable generation and audio-driven animation.
@@ 3 Dataset
We build our animation dataset according to the observation that high quality text-video pairs are the cornerstone of video generation, which is proved by recent researches Polyak et al. (2024). In this section, we give a detailed description of our animation dataset and the evaluation benchmark.
Animation Dataset Construction: We build a pipeline to get high-quality text-video pairs among $1$ million raw animation videos. First of all, we use scene detectionBreakthrough (2024) to divide raw videos into clips. Then, for each video clip, we construct a filter rule from four dimensions: text-cover region, optical flow score, aesthetic score, and number of frames. The filter rule is gradually built up through the observations in model training. In detail, the text-cover region scoreBaek et al. (2019) can drop those clips with text overlay similar to end credits. Optical flow score princeton vl (2020) prevents those clips with still images or quick flashback scenes. Aesthetic score christophschuhmann (2022) is utilized to preserve clips with high artistic quality. Besides, we retain the video clips whose duration is among $2s$-$20s$ according to the number of frames. Furthermore, we collected 0.5M high-quality animation videos along with their corresponding captions to create video-text pairs, which were used to fine-tune Qwen-VL2Wang et al. (2024). After fine-tuning, the model provides more accurate descriptions of characters, scenes, and action details in animation content. After the steps mentioned above, about $10\%$ clips (more than 10M clips) with captions can be retained in the training step.
In addition, since occupationally-generated animation videos typically have significantly higher production costs and quality compared to user-generated animation content, we fine-tuned Qwen-VL2 based on these data. This model is then utilized to filter higher-quality clips to further improve the model’s performance. Specifically, during the training process, we adjust the proportions of specific training data (e.g., talking and motion amplitude) according to the observed performance.
Benchmark Dataset Construction: Since there is currently no benchmark dataset specifically designed for animation content, we construct a benchmark dataset manually to compare the generation videos between our model and other recent researches. $948$ animation clips are collected and labeled with different actions, e.g., talking, walking, running, eating, and so on. Among them, there are $857$ 2D animation clips and $91$ 3D clips. These action labels are summarized from more than $100$ common actions with human annotation. Each label contains $10$-$30$ video clips. The corresponding text prompt is generated by fine-tuned Qwen-VL2 (mentioned above) at first, then is corrected manually to guarantee the text-video alignment. (More details in section 5 and 6)
@@ 4 Method
In this section, we present an effective approach for animation video generation using a diffusion transformer architecture. Section 4.1 provides an overview of the foundational diffusion transformer model. In section 4.2, we introduce a spatiotemporal mask module that extends the model, enabling crucial animation production functions such as image-to-video generation, frame interpolation, and localized image-guided animation within a unified framework. These enhancements are essential for professional animation production. Finally, section 4.3 details the supervised fine-tuning strategy employed on the animation dataset.
@@ 4.1 DIT-based Video Generation Model
We adopt a DiT-basedPeebles and Xie (2023) text-to-video diffusion model as the foundation model. As shown in Fig. 3, the model leverages the three components to achieve coherent, high-resolution videos aligned with text prompts.
Figure 3: Method. This figure illustrates the Masked Diffusion Transformer framework for animation video generation, designed to support various spatiotemporal conditioning methods for precise and flexible animation control. A 3D Causal VAE compresses spatial-temporal features into a latent representation, generating the guide feature sequence $G$, while a reprojection network constructs the mask sequence $M$. These components, combined with noise and prompt’s feature, serve as input to the Diffusion Transformer. The transformer employs techniques such as patchify, 3D-RoPE embeddings, and 3D full attention to effectively capture spatial-temporal dependencies. This framework seamlessly integrates keyframe interpolation, motion control, and mid-frame extension, simplifying animation production and enhancing creative possibilities.
3D Casual VAE used in video generation frameworksGupta et al. (2023); Yu et al. (2023)serves as a specialized encoder-decoder architecture tailored for spatiotemporal data compression. This 3D VAE compresses videos across both spatial and temporal dimensions, significantly reducing the diffusion model computing. We follow the approach of Yang et al. Yang et al. (2024) to extract latent features, transforming the original video with dimensions $(W,H,T,3)$ into a latent representation of shape $(W/8,H/8,T/4,16)$.
Patchify is a critical step for adapting vision tasks to transformer-based architectures Alexey (2020). Given an input video of size $T\times H\times W\times C$, it is split spatio into patches of size $P\times P$, and temporal into size $\ Q$ resulting in $(T/Q)\times(H/P)\times(W/P)\times C$ patches. This method enables efficient high-dimensional data processing by reducing complexity while retaining local spatial information.
3D Full Attention We propose 3D full attention module for spatial and temporal modeling considering the great success of long-context training in LLMDubey et al. (2024) and foundation video generation modelYang et al. (2024); Polyak et al. (2024).
@@
To stabilize training, we use the v-prediction loss Salimans and Ho (2022), where $v=\sqrt{1-\alpha_{t}}\,x_{0}-\sqrt{\alpha_{t}}\,\epsilon$ and the loss becomes
@@ 4.2 Spatiotemporal Condition Model
Motion Control Our framework enables precise control over motion regions addressing the limitations of text-based control in these aspects. This approach enhances artists’ control over video content, allowing them to express their creativity while significantly reducing their workload.
@@ 4.2.2 Motion Area Condition
This framework can also support spatial motion area conditions inspired by Dai et.alDai et al. (2023). Given the image condition $F_{p_{i}}$, and motion area condition is represented by mask $M_{F}$, the same shape with $F_{p_{i}}$. Motion area in $M_{F}$ is labeled 1, other place is set to 0. As equation 1 in 4.2.1, for guide frame position $p_{i}$, set $M_{p_{i}}=M_{F}$. The data processing and training pipeline can be summarized as follows: Constructing video-mask pairs, we first construct paired training data consisting of videos and corresponding masks. Using a foreground detector by Kim et.al Kim et al. (2022), we detect the foreground region in the first frame of the video. This region is then tracked across subsequent frames to generate a foreground mask for each frame. Union of foreground masks, the per-frame foreground masks are combined to create a unified mask $M_{F}$, representing the union of all foreground regions across the video. Video latent post-processing, for the video latent representation $z_{0}$, non-moving regions are set to the latent features of the guide image, ensuring static areas adhere to the guide. LoRA-based conditional training, we train the conditional guidance model using Low-Rank Adaptation (LoRA) with a parameter size of 0.27B. This approach significantly reduces computational requirements while enabling efficient model training.
@@ 4.3 Supervised Fine-Tuning
Multi-Task Learning Compared to the physically consistent motion patterns in the real world, animation styles, and motion dynamics can vary significantly across different works. This domain gap between datasets often leads to substantial quality differences in videos generated from guide frames with different artistic styles. We incorporate image generation into a multi-task training framework to improve the model’s generalization across diverse art styles. Experimental results in the appendix demonstrate that this approach effectively reduces the quality gap in video generation caused by stylistic differences in guide frames.
@@
Mask Strategy During training, we unmask the first, last, and other frames obtained through uniform sampling with a 50% probability. This strategy equips the model with the ability to handle arbitrary guidance, enabling it to perform tasks such as in-betweening, first-frame continuation, and arbitrary frame guidance, as discussed in Section 4.2.1.
In practice, we also employed several other effective training strategies, such as weak to strong training, generated subtitle removal, and temporal multi-resolution training. Detailed training procedures can be found in the appendix.
5 Benchmark
To evaluate the effects of the animation video generation models, we build a comprehensive benchmark dataset, as mentioned in section 3. To give a fair comparison of different methods, we define $6$ basic dimensions to describe the quality of the generated videos. Then, we introduce the human annotation and evaluation metrics, which are partly based on the annotation results.
5.1 Evaluation Dimensions
In fact, the essential concepts for evaluating a animation video generation model are visual appearance and visual consistency. Visual appearance describes the basic quality, which is only concerned with the generation of videos themselves, including visual smoothness, visual motion, and visual appeal, while visual consistency considers the text-video, image-video, and character consistency, respectively.
Inspired by VbenchHuang et al. (2024), we first adopted a similar approach to evaluate visual appearance and consistency, such as calculating the clip score between frames to analyze distortions and using aesthetic scores to estimate aesthetic quality. However, we found that these scores often showed significant discrepancies from subjective human experiences on animation data and there was no clear distinction between different methods. As shown in Fig.4, the bottom video received a lower score due to its larger motion amplitude. However, the top video contains visible distortions that are easily noticeable to humans, yet these are not reflected in the scoring metrics.
Therefore, we adopted a regression-based approach for certain dimensions to learn human scoring standards. A detailed description of six evaluated metrics is given as follows. These evaluation criteria are specifically designed for animation data and align closely with human subjective experience.
Figure 4: Video generated by Opensora-V1.2 (top) and Vidu (bottom). The top video received a higher score despite containing noticeable distortions.
5.2 Human Annotation
In the experiment, each case in the benchmark has $6$ generation video clips with different participant models. $20$ expert volunteers give their rating (1-5, 5 is the best) from above $6$ dimensions, the detail scores are shown in Section 6.1.
5.3 Visual Appearance
We evaluate the basic quality of a video clip from three aspects: visual smoothness, visual motion, and visual appeal. We constructed an evaluation model training set to develop an evaluation model aligned with human scoring. We generated $5000$ animated video clips using various models described in Section 6.1. These samples were manually annotated following the approach outlined in Section 5.2, serving as the ground truth for training the evaluation model.
Visual Smoothness Our goal was to learn the human standards to evaluate video smoothness and to be able to identify animated videos with distortions. Thus, we trained a model to regress human-generated scores to avoid the impact of motion on visual content. The training set consists of generated clips and annotations, mentioned in Section 5.3. To enhance the robustness of the model, we also incorporated hundreds of anime videos as the highest-scoring examples.
Visual Motion We employ a model based on the ActionCLIP Wang et al. (2021) framework to train a motion-scoring model that evaluates the magnitude of the primary motion in animation videos. About $2$ thousand animation video clips and their corresponding motion captions are collected into $6$ degrees of movement amplitude (from stillness to significant motion) to finetune the action model. Finally, the motion score is obtained from the similarity score between the designed motion prompt and the participant video.
$S_{motion}=Cos(MCLIP(V),MCLIP(T_{m})),$ (2)
where $MCLIP$ denotes the finetuning action model. $V$ represents the generation video and $T_{m}$ denotes the designed motion prompt. (Details is provided in supplementary)
Visual Appeal We define visual appeal score to reflect the basic effects of video generation. As discussed in Section 5.1, the aesthetic score models used in previous studies were trained on real-world datasets such as LAION-5B, simply calculating aesthetic scores on animation data does not provide sufficient differentiation, as the results from various methods are indistinguishable.
Using the key frame extraction method to collect the key frames in the video first, then we train a regress model on the evaluation model training set to learn human aesthetic standards. The formulation shows as follows:
$S_{appeal}=Aes(SigLIP(I_{0,1,\dots,K}))\quad I_{i}\in{KeyFrm(V)},$ (3)
where $KeyFrm$, $SigLIP$ and $Aes$ denote the key frame extraction method, feature encoder method and aesthetic evaluation method, and $K$ denotes the number of the keyframes.
5.4 Visual Consistency
Three factors are considered to evaluate the visual consistency of the generation video: text-video, image-video, and character consistency, respectively.
Text-video Consistency To evaluate the text-video consistency, we finetune the vision encoder and the text encoder modules with a regression head according to animation video-text pairs according to the training set in Section 5.3. The formulation is shown as follows:
$S_{tvc}=Reg(E_{v},E_{t}),$ (4)
where $Reg$ denotes the regression head, and $E_{v}$, $E_{t}$ denote the vision and text encoder.
Image-video Consistency In the I2V situation, the participant image, as an input factor, should ensure that its style is consistent with the generated videos. Similar to text-video consistency, we combine a vision encoder with a regression head to evaluate the score. The model is also fine-tuned on the training set in Section 5.3. The formulation lists as follows:
$S_{ivc}=Reg(E_{v}(V),E_{v}(I_{p})),$ (5)
where $V$ and $I_{p}$ denote the participant video clip and the input image.
Character Consistency Character consistency is a crucial factor in animation video generation. When the character generated by the protagonist in the animation changes, even if the quality of the video is great, it still has the risk of infringement. Hence, we design a set of procedures including detection, segmentation, and recognition. We apply GroudingDino Ren et al. (2024) and SAM Ravi et al. (2024) to achieve character mask extraction for each frame in the videos. Then, we finetune a BLIP-based model Li et al. (2022) to establish connections between each mask and the animation IP character. In detail, thousands of source video clips with their characters labeled are treated as training sets to obtain and store the reliable features from BLIP-based model. In the evaluation step, we get the score of character consistency by calculating the cosine similarity between the generated and stored character’s features.
$S_{IPc}=\frac{1}{S}\sum^{S}_{i}Cos(BLIP(M_{i}),fea_{c}),$ (6)
where $S$ denotes the number of sample frames, $M_{i}$ denotes the mask obtained from GroudingDino and SAM methods, and $fea_{c}$ denotes the stored character’s features.